=== 2019/1/22(Tue.) ===

=== Weekend ===


==== Transition from morning work (*:responsible person) ====
 * Haibo starts joining commissioning work. He will stay 20days.

==== Report of work (*:responsible person) ====

 * BS: Two cables were not connected well, so actuation was not correct when measuring TF day before yesterday.
  * TF was measured yesterday but still TF showed something different from the last correct measurement.
  * FR moved 800um.
  * Anyway alignment can be proceeded probably this afternoon.

 * CRY: we found disconnected cables at ITMX.

 * OMC: OMMT was aligned, and it was found that OSEM was touched somewhere.
 * PEM: Ondotori is still not working because of software issue. 


==== Work plan (*:responsible person) ====

 * COMM: we will check PR3 first, then SR3.
  * Access to PR3 is quite difficult, gate valve needs to be removed or optical table needs to be fix.
  * we will decide whether we need clean the mirror surface or not.
  * Once mirror inspection and cleaning (if necessary), then will start alignment for REFL using reflected beam from ITMX.
 * COMM: PDA will be placed EYA.

 * SR3, SR2: height is OK. FR will be tuned.
  * SR3 surface will be inspected.